OpenStudy (anonymous):

Some one explain please For the validity of a binomial expansion[ (1+x)^n], some books say its valid if |x|>1 and some books say valid if -1>x>1 what do they both mean and are they the same?

OpenStudy (jdoe0001):

well, the binomial theorem works with exponents that are integer now if you have a negative exponent that's an integer, well \(\bf a^{-n} = \cfrac{1}{a^n}\qquad \qquad (x+y)^{-n} = \cfrac{1}{(x+y)^{n}}\)

OpenStudy (anonymous):

Yeah but what does -1>x>1 mean in math notation?

OpenStudy (anonymous):

And does |x|>1 mean only fractions?

zepdrix (zepdrix):

|x|>1 and -1>x>1 are the same thing :)

OpenStudy (jdoe0001):

well, -1 > x # "x" is less than -1 keep in mind that on the negative values, the farther from the 0, the smaller, thus -25 is smaller than -1 -1 > x # will mean \(\bf (-\infty, 1)\)

OpenStudy (anonymous):

@Zepdrix But -0.5 can fall in range range of -1>x>1 but not |x|>1

OpenStudy (jdoe0001):

but -0.5 is not an integer

OpenStudy (jdoe0001):

though ... -0.5 is greater than -1, and outside the range indicated

OpenStudy (anonymous):

Sorry i've been using the wrong sign I mean |x|<1 and -1<x<1

zepdrix (zepdrix):

Oh I'm sorry, I misread what your wrote. I meant to say, -1>x>1 is equivalent to |x|<1 Not greater than.

zepdrix (zepdrix):

Oh lol

OpenStudy (jdoe0001):

hmmm, actually I have it right, even with the > lhehehhe

OpenStudy (jdoe0001):

well, -1 < x # "x" is greater than -1 keep in mind that on the negative values, the farther from the 0, the smaller, thus -25 is smaller than -1 -1 < x # will mean \(\bf (-\infty, 1)\)

OpenStudy (jdoe0001):

hmmm, wait.... a sec

OpenStudy (jdoe0001):

any greater .... will be to the rigtht... ok ... well have, better read closely

OpenStudy (jdoe0001):

hmm |x|<1 and -1<x<1 that only leaves room for rationals

OpenStudy (jdoe0001):

maybe they mean \(\bf a^{-n} = \cfrac{1}{a^n}\qquad (x+y)^{-n} = \cfrac{1}{(x+y)^{n}} \qquad \large (x+y)^{\frac{a}{b}} = \sqrt[b]{(x+y)^a}\)

OpenStudy (anonymous):

Yes yes but why does -0.5 fall in range of -1<x<1 but not |x|<1

OpenStudy (anonymous):

They mean when n is not a positive integer

OpenStudy (anonymous):

@jdoe0001

zepdrix (zepdrix):

-0.5 DOES fall in the range of |x|<1 It didn't when you had the equality sign flipped, but it does now :O

zepdrix (zepdrix):

\[\Large |-0.5|<1 \qquad\to\qquad 0.5<1\]

OpenStudy (anonymous):

Ohh yeeeeeaaaahh!! Sorry

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

Thanks you both, can't give two medals so I won't give any X_X

zepdrix (zepdrix):

I have too many medals XD you can give it to jdoe if you want. hehe

OpenStudy (anonymous):

Can I have some help here

OpenStudy (anonymous):

Letf(x)= 9x2+4 . (2x + 1)(x − 2)2 (i) Express f(x) in partial fractions. (ii) Show that, when x is sufficiently small for x3 and higher powers to be neglected, f(x) = 1 − x + 5x2. What does it mean to show that x is sufficiently small for x^3 ?